OpenStudy (anonymous):

2+2log(4)Y = 3log(2)X

OpenStudy (anonymous):

\[2+2\log _{4}y = 3\log _{2}x \]

OpenStudy (anonymous):

how????????

OpenStudy (anonymous):

... \[\LARGE 2+2\cdot \log_4(y)=3\cdot \log_{2}(x)\] \[\LARGE 2+2\cdot \log_{2^2}(y)=3\cdot \log_{2}(x)\] \[\LARGE 2+2\cdot\frac12 \cdot \log_{2}(y)=3\cdot \log_{2}(x)\] \[\LARGE 2+ \log_{2}(y)= \log_{2}(x^3)\] \[\LARGE 2 = \log_{2}(x^3)-\log_{2}(y)\] \[\LARGE 2 = \log_{2}\left(\frac{x^3}{y}\right)\] \[\LARGE 2^2 = \frac{x^3}{y} \] now solve for each one you like...

OpenStudy (anonymous):

how you guys get so smart lol

OpenStudy (anonymous):

You'll be smarter when you'll reach "our" age. ;)